When an object is released and falls freely, its height , h metres, above the ground after t seconds is represented by the relation h = -0.5t2g + d , where g is the acceleration due to gravity and d is the initial height of the object before it is released. Since the strength of gravity varies from planet to planet, the value of g is specific to each planet.
Suppose a rock is dropped from a height of 100 m on Mars and also on Venus.
On Mars, g= 3.7m/s2 and on Venus, g= 8.9 m/s2. How long does it take the rock to fall to a height of 25 m on each planet ?

Answer by Alan3354(69443)

When an object is released and falls freely, its height , h metres, above the ground after t seconds is represented by the relation h = -0.5t2g + d , where g is the acceleration due to gravity and d is the initial height of the object before it is released. Since the strength of gravity varies from planet to planet, the value of g is specific to each planet.
Suppose a rock is dropped from a height of 100 m on Mars and also on Venus.
On Mars, g= 3.7m/s2 and on Venus, g= 8.9 m/s2. How long does it take the rock to fall to a height of 25 m on each planet ?
==============
On Mars, g= 3.7m/s2
h(t) = -1.85t^2 + 100 = 25
t^2 = 75/1.85
t =~ 6.367 seconds
-------
Do Venus the same way.